# In-medium changes of nucleon cross sections tested in neutrino-induced reactions

### v1

@@ Acknowledgements.
- [1] B. Ter Haar and R. Malfliet, Phys. Lett. B 172, 10-16 (1986) doi:10.1016/0370-2693(86)90207-8
@@
- [4] M. Henri, O. Lopez, D. Durand, B. Borderie, R. Bougault, A. Chbihi, Q. Fable, J. D. Frankland, E. Galichet and D. Gruyer, et al. Phys. Rev. C 101, no.6, 064622 (2020) doi:10.1103/PhysRevC.101.064622
@@
- [9] A.A. Aguilar-Arevalo et al. [MiniBooNE], Phys. Rev. D 79, 072002 (2009) doi:10.1103/PhysRevD.79.072002 [arXiv:0806.1449 [hep-ex]]
@@
- [12] T. Song and C. M. Ko, Phys. Rev. C 91, no.1, 014901 (2015) doi:10.1103/PhysRevC.91.014901
- [13] E. Oset and L. L. Salcedo, Nucl. Phys. A 468, 631-652 (1987) doi:10.1016/0375-9474(87)90185-0
- [14] P. Abratenko et al. [MicroBooNE], [[arXiv:2402.19216 [hep-ex]].
@@
- [20] T. E. O. Ericson and W. Weise, Clarendon Press, 1988, ISBN 978-0-19-852008-5
- [21] S. Gardiner, [arXiv:2401.04065 [hep-ex]].

### v3

@@ Acknowledgements.
- [1] B. Ter Haar and R. Malfliet, Phys. Lett. B 172, 10-16 (1986) doi:10.1016/0370-2693(86)90207-8.
@@
- [4] M. Henri, O. Lopez, D. Durand, B. Borderie, R. Bougault, A. Chbihi, Q. Fable, J. D. Frankland, E. Galichet and D. Gruyer, et al. Phys. Rev. C 101, no.6, 064622 (2020) doi:10.1103/PhysRevC.101.064622.
@@
- [7] Jiang, L. et al. [Jefferson Lab Hall A Phys. Rev. D 105, 112002 (2022) doi:10.1103/PhysRevD.105.112002 [arXiv:2203.01748 [nucl-ex]].
- [8] L. Gu et al. [Jefferson Lab Hall A], Phys. Rev. C 103, 034604 (2021) doi:10.1103/PhysRevC.103.034604 [arXiv:2012.11466 [nucl-ex]].
- [9] M. Duer et al. [CLAS], Nature 560, 617–621 (2018) doi:10.1038/s41586-018-0400-z
- [10] C. Barbieri, D. Rohe, I. Sick, and Louk Lapikas, Phys. Lett. B 608, 47-52 (2005) doi:10.1016/j.physletb.2004.12.072 [arXiv:nucl-th/0411066].
@@
- [13] A.A. Aguilar-Arevalo et al. [MiniBooNE], Phys. Rev. D 79, 072002 (2009) doi:10.1103/PhysRevD.79.072002 [arXiv:0806.1449 [hep-ex]].
@@
- [16] T. Song and C. M. Ko, Phys. Rev. C 91, no.1, 014901 (2015) doi:10.1103/PhysRevC.91.014901.
- [17] E. Oset and L. L. Salcedo, Nucl. Phys. A 468, 631-652 (1987) doi:10.1016/0375-9474(87)90185-0.
- [18] R. González-Jiménez, K. Niewczas and N. Jachowicz, Phys. Rev. D 97, no.1, 013004 (2018) doi:10.1103/PhysRevD.97.013004 [arXiv:1710.08374 [nucl-th]].
- [19] P. Abratenko et al. [MicroBooNE], Phys. Rev. D 110, 013006 (2024) doi:10.1103/PhysRevD.110.013006 [[arXiv:2402.19216 [hep-ex]].
@@
- [33] P. Abratenko et al. [MicroBooNE], [arXiv:2406.10583 [hep-ex]].
- [34] Airapetian, A. et al. [HERMES], Phys. Rev. Lett. 96, 162301 (2006) [arXiv:hep-ex/0510030 [hep-ex]].